# Improving Group Robustness on Spurious Correlation Requires Preciser Group Inference

### v1

@@ Abstract
Standard empirical risk minimization (ERM) models may prioritize learning spurious correlations between spurious features and true labels, leading to poor accuracy on groups where these correlations do not hold. Mitigating this issue often requires expensive spurious attribute (group) labels or relies on trained ERM models to infer group labels when group information is unavailable. However, the significant performance gap in worst-group accuracy between using pseudo group labels and using oracle group labels inspires us to consider further improving group robustness through preciser group inference. Therefore, we propose GIC, a novel method that accurately infers group labels, resulting in improved worst-group performance. GIC trains a spurious attribute classifier based on two key properties of spurious correlations: (1) high correlation between spurious attributes and true labels, and (2) variability in this correlation between datasets with different group distributions. Empirical studies on multiple datasets demonstrate the effectiveness of GIC in inferring group labels, and combining GIC with various downstream invariant learning methods improves worst-group accuracy, showcasing its powerful flexibility. Additionally, through analyzing the misclassifications in GIC, we identify an interesting phenomenon called semantic consistency, which may contribute to better decoupling the association between spurious attributes and labels, thereby mitigating spurious correlation.
@@ 1 Introduction
The presence of spurious correlation causes machine learning models to fail on certain groups of samples, even when achieving high accuracy on average group [38, 4, 20, 13]. Deep networks trained via standard empirical risk minimization (ERM) may be biased by the spurious attributes/features such as the image backgrounds [4, 16], which admit different correlations with the true labels for the different groups of data. As a result, the model typically performs well on the majority group of data that aligns with the spurious attributes but performs poorly on certain groups of data that have no or even opposite correlations with spurious attributes. This critical issue is widespread in many fields, such as medical AI, facial recognition and sentiment analysis [7, 46, 19, 3, 40].
@@
There have emerged a series of previous works that aim to develop more robust training methods to improve the worst-group performance when group information is available. For instance, one can focus on particularly training over the worst-group risk [40] via group distributional robust optimization (Group DRO). Additionally, the spurious correlation can also be eliminated by balancing the risk in different groups, which can be achieved via feature reweighting [22] or selective Mixup [50]. These methods have been demonstrated to be effective to mitigate the spurious correlations, leading to substantial improvements over ERM.
@@
However, group labeling is expensive, labor-intensive, and human-biased. It is more important and practical to develop robust algorithms for improving worst-group accuracy when the group information is unavailable [31, 51]. In such a scenario, a line of recent works [43, 34, 1, 31, 51, 10] propose to first infer the data groups, and then seek to train a robust model according to the estimated group information. For instance, JTT [31] identifies the data being misclassified by the ERM as the minority group, and trains a robust model via upweighting the examples in the minority group; CnC [51] infers pseudo group labels via ERM and develops a contrastive learning method to train the final model.
@@
The primary focus of the aforementioned research is on invariant learning, i.e., leveraging the group information obtained from ERM to learn the invariant attributes. However, the accuracy of group label inference has been overlooked, making the existing group inference-based methods perform significantly worse than those using oracle group labels. Very recently, some group inference methods have been developed as alternatives to ERM, such as EIIL [12], SSA [35], DISC [48], ZIN [30]. However, they may either struggle with complex spurious correlations (e.g., EIIL) [30], or require additional information (e.g., human inspection, few spurious annotations, etc) that is related to the group labels (e.g., SSA, DISC, and ZIN). Therefore, noticing the performance shortcomings and applicability limitations in existing group inference methods, we raise an important yet challenging question: can we develop a more accurate group inference method to mitigate spurious correlations without relying on any additional information?
@@
In this work, we propose GIC (group inference via data comparison), a novel, practical, and accurate method to infer group labels, which can be then seamlessly incorporated into a variety of oracle group label-based robust learning methods to improve the worst-group performance. In particular, GIC adopts a (unlabeled) comparison dataset, which has (slightly) different group distribution compared to the training dataset, and infers the group information in a contrastive manner. More specifically, GIC trains an alternative classifier to assign spurious attribute labels to data points solely on their spurious attributes, where the training objective is designed to encourage discovering the distribution discrepancy between the comparison and training dataset while maintaining the high correlation between the spurious attribute label and true label in the training dataset. Notably, unlike previous methods that may rely on (partial) group labels or human inspections, obtaining the comparison data does not necessitate any assumptions. It does not need to be labeled and can be obtained from various sources, such as a validation set with true validation labels, a completely unlabeled subset sampled from the test data, or even directly resampled from the training dataset in a non-uniform manner. We highlight the main contributions as follows:
@@
    We show that the proposed GIC can seamlessly integrate with multiple invariant learning algorithms to improve the worst-group accuracy. In Section 4.3 and 4.4, GIC is successfully combined with Mixup [50], GroupDRO [12], Upsample [31], and Subsample [22]. It can be seen that GIC consistently outperform baselines in terms of the worst-group accuracy, for all candidate invariant learning algorithms. More importantly, when integrated with Mixup, our model improves over the state-of-the-art for nearly all tasks when the group label is unavaliable. Additionally, the average and worst-group accuracy of our model can almost match that of using oracle group labels directly. This further justifies the effectiveness of our group inference method.
@@ 2 Related Work
Recent research show that the traditional ERM learns both spurious and invariant features [22, 21, 39, 10]. However, with strong spurious correlations, ERM tends to prioritize learning the spurious features [10], which hinders its ability to generalize on data where these spurious correlations are absent. To tackle this issue, invariant feature learning, referred to as invariant learning for simplicity, has been introduced to learn invariant representations. One notable approach is group robustness methods.
When group labels are available, some classical group robustness methods, such as GroupDRO [40], attempt to minimize the worst-group loss instead of the average loss using oracle group labels. Other methods aim to achieve invariant learning by balancing the majority and minority groups, such as reweighting [41], regularization [9], and downsampling [22]. Additionally, approaches like semi-supervised learning [35, 44] or Mixup [50], which selectively combine samples with matching labels but differing spurious attributes or matching spurious attributes but differing labels, are also be considered to improve the the worst-group accuracy. Some methods attempt to infer group labels by training a simple ERM [7, 46, 19, 3, 40], maximizing the GroupDRO loss [12], or even introducing human knowledge [30, 48] when group information is unkonwn. However, these inferred group methods have performance gaps compared to group annotation utilized methods and may not be applicable when prior information is unavailable. In this work, we also focus on group robustness without any group information.
We point out that, although we aims to learn invariant features to train more robust and generalizable models, similar to classical domain adaptation [6, 33] methods, such as UDA [14] and DANN [15], our setting differs from them which divide the training data into source and target domains and use the source and target features, along with the source labels, to transfer knowledge to the specific target domain [51]. In contrast, we do not have a natural source and target domains and strive to enhance the accuracy of specific groups affected by spurious correlations, without prior knowledge of training sample domains or spurious attributes.
@@ 3.1 Problem Setup
Consider the training dataset $\mathcal{D}$, which comprises $n$ data point-label pairs, where each data point-label pair belongs to some group $g\in\mathcal{G}$, denoted as $\mathcal{D}=\{(\mathbf{x}_{i},y_{i},g_{i})\}_{i=1}^{n}$. Following previous work [31, 51, 50, 49], we consider each group $g=(y,a)$ to be jointly defined by the label $y\in\mathcal{Y}$ and unobserved spurious attributes $a\in\mathcal{A}$ (i.e., $\mathcal{G}=\mathcal{Y}\times\mathcal{A}$), where the spurious attribute $a$ is spuriously correlated with the label $y$ (e.g., the image background and the image label). We denote $y_{s}$ as the spurious attribute label of $a$ (e.g., the label of the image background, which can be “land” vs “water” or “red” vs “green”). The groups $g$ in dataset $\mathcal{D}$ often encounter the imbalance issue, where ERM training tends to mostly depend on majority groups, thus may memorize the spurious correlation contributed by these groups. However, such spurious correlations are often absent or even oppositely appearing in minority groups. Therefore, our ultimate objective is to train a robust model $f_{\mathrm{robust}}(\theta)$, parameterized by $\theta\in{\Theta}$, capable of classifying an input $\mathbf{x}_{i}$ to a label $y_{i}$ regardless of whether it belongs to the majority or minority group. This is referred to as the worst-group accuracy defined as follows:
${\max_{g\in\mathcal{G}}\mathbb{E}[\mathbbm{1}[f_{\theta}(\mathbf{x})\neq y]|g]},$ (1)
@@ 3.2 GIC: Group Inference via data Comparison
$\hat{y}_{s}=f_{\mathrm{GIC}}(\mathbf{z};\mathbf{w}),$ (3)
where $\mathbf{w}$ is weights of GIC model $f_{\mathrm{GIC}}$. Then using the predicted $\hat{y}_{s}$, GIC partitions the data into different groups, i.e., $\hat{g}=(y,\hat{y}_{s})$. The inferred group $\hat{g}$ can be integrated into downstream invariant learning methods, aiding in training the robust model $f_{\mathrm{robust}}$.
Comparison data. Before formally introducing the algorithm details of GIC, we first introduce the comparison data. We assume access to a dataset $\mathcal{C}$, where its group distribution (slightly) differs from the training data $\mathcal{D}$, e.g., the group distribution of $\mathcal{D}$ is $(g_{1},g_{2})=(0.1,0.9)$, while the group distribution of $\mathcal{C}$ is $(g_{1},g_{2})=(0.2,0.8)$. We refer to $\mathcal{C}$ as comparison data, which can have true labels or be unlabeled. We remark that the comparison data is easy to obtain, which will be thoroughly discussed in Section 3.4.
Based on the definition of spurious attribute labels $y_{s}$, an ideal spurious attribute prediction $\hat{y}_{s}$ should exhibit the following two properties: (1) It should be highly correlated with the true label $y$. It is this high correlation between the spurious attribute label and the true label that biases the ERM-based model training. (2) The correlation between $\hat{y}_{s}$ and $y$ varies across different datasets. Unlike the invariant attribute label, which is always equivalent to the true label, the correlation between the spurious attribute label and the true label is spurious, changing with the spurious attribute distribution across different datasets. Consider an extreme case, the spurious background attribute for both waterbirds and landbirds in the comparison data accounts for 50%, while in the training set, the spurious attribute label is completely equivalent to the true label (e.g., $y_{s}=y$). The equivalence correlation that exists in the training data does not hold in the comparison data.
Based on the mentioned two properties, we design the optimization objective of GIC, which consists of two terms: (1) Correlation Term, used to describe the high correlation between $\hat{y}_{s}$ and $y$ in the training data; (2) Spurious Term, used to describe the discrepancy in the correlation between the training and comparison data, aiming to emphasize the correlation is spurious rather than invariant. By jointly optimizing the above objectives, GIC is encouraged to train a spurious attribute classifier that yields the spurious attribute prediction $\hat{y}_{s}$, as shown in Equation (3).
Correlation Term. To encourage the high correlation between $y$ and $\hat{y}_{s}$ in the training set, we consider the following optimization objective:
@@
where $\hat{y}^{tr}_{s,\mathbf{w}}=f_{\mathrm{GIC}}(\mathbf{z}^{tr};\mathbf{w})$ is the predicted spurious attribute label on training data. Here, mutual information is considered due to its widespread usage to measure the correlation or dependency between random variables [29, 24, 36, 45]. We aim to maximize the mutual information between $y_{s}$ and $y$, encouraging the predicted spurious attribute label $y_{s}$ from GIC to be highly correlated with the true label $y$.
Solely satisfying Equation (4) is not enough as invariant attribute information contained in feature representations [22, 10] can also exhibit high correlation with true labels. To aid GIC in learning spurious attributes instead of invariant ones, we introduce an additional spurious term.
Spurious Term. We use conditional probability to describe the correlation between $\hat{y}_{s}$ and $y$, i.e., $\mathbb{P}(y|\hat{y}_{s})$. Then let $\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s})$ and $\mathbb{P}(y^{c}|\hat{y}^{c}_{s})$ be the conditional distributions in the training and comparison dataset respectively, we will characterize their discrepancy. Since we are comparing distributions with the same support, we opt for the simplest and widely used choice [1, 27], the KL-divergence. Then, we maximize the following objective to encourage GIC to learn spurious attributes:
@@
The essence of maximizing Equation (5) is to encourage GIC to learn spurious attributes by violating the invariant learning principle [12]. If $\hat{y}_{s}$ is inferred based on invariant attributes rather than spurious ones, then $\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})$ and $\mathbb{P}(y^{c}|\hat{y}^{c}_{s,\mathbf{w}})$ will be identical and then
@@
Conversely, if $\hat{y}_{s}$ is based on spurious attributes rather than invariant attributes, the inequality
@@
will hold, where the equality holds true only when the training data and the comparison data have the same group distribution. Therefore, maximizing Equation (5) entails encouraging GIC to prioritize learning spurious attributes over invariant attributes.
@@
Obtain $\mathbf{z}^{tr}=\Phi(\mathbf{x^{tr}}),\mathbf{z}^{c}=\Phi(\mathbf{x}^{c})$ where $\mathbf{x^{c}}\in\mathcal{D},\mathbf{x}^{c}\in\mathcal{C}$.
@@
Initialize the spurious attribute classifier $f_{\mathrm{GIC}}$.
1:for epoch $1$ to $K$ do
2:     if the true label of $\mathcal{C}$ is available then
3:         Optimizing Equation (10) to update $f_{\mathrm{GIC}}$.
4:     else
5:         Optimizing Equation (11) to update $f_{\mathrm{GIC}}$.
6:     end if
7:end for
8:Infer spurious attribute labels $\hat{y}^{tr}_{s}=f_{\mathrm{GIC}}(\mathbf{z}^{tr})$.
9:return group labels $\hat{g}=(y^{tr},\hat{y}^{tr}_{s})$.
Overall Objective. By combining (4) and (5), we derive the overall objective of GIC as follows:
@@
where $\gamma>0$ is a weighting parameter used to balance Correlation Term and Spurious Term.
However, the overall objective of GIC faces certain practical issues. Firstly, the mutual information term is difficult to accurately estimate [37, 5]. Secondly, Equation (8) cannot handle the situation where the comparison data is unlabeled, which limits the applicability of GIC in various scenarios.
We first replace the mutual information $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$ with the cross-entropy $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$ to achieve accurate estimation. A detailed proof in Appendix A.1 demonstrates that $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$ is, in fact, a lower bound of $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$. Therefore, maximizing $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$ can be approximated by maximizing $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$. This replacement aligns with intuition because maximizing mutual information between $y^{tr}$ and $\hat{y}^{tr}_{s,\mathbf{w}}$ essentially encourages a closer alignment of their distributions, which is consistent with the objective of minimizing $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$.
We then extend GIC to the case where true labels $y^{c}$ of the comparison data are not available. We substitute the true labels $y^{c}$ with the comparison data’s feature representation $\mathbf{z}^{c}$, which strongly associates with $y^{c}$ and is always accessible. We present the following theorem:

### v2

@@ Abstract
Standard empirical risk minimization (ERM) models may prioritize learning spurious correlations between spurious features and true labels, leading to poor accuracy on groups where these correlations do not hold. Mitigating this issue often requires expensive spurious attribute (group) labels or relies on trained ERM models to infer group labels when group information is unavailable. However, the significant performance gap in worst-group accuracy between using pseudo group labels and using oracle group labels inspires us to consider further improving group robustness through preciser group inference. Therefore, we propose GIC, a novel method that accurately infers group labels, resulting in improved worst-group performance. GIC trains a spurious attribute classifier based on two key properties of spurious correlations: (1) high correlation between spurious attributes and true labels, and (2) variability in this correlation between datasets with different group distributions. Empirical studies on multiple datasets demonstrate the effectiveness of GIC in inferring group labels, and combining GIC with various downstream invariant learning methods improves worst-group accuracy, showcasing its powerful flexibility. Additionally, through analyzing the misclassifications in GIC, we identify an interesting phenomenon called semantic consistency, which may contribute to better decoupling the association between spurious attributes and labels, thereby mitigating spurious correlation. The code for GIC is available at https://github.com/yujinhanml/GIC.
Machine Learning, ICML
@@ 1 Introduction
The presence of spurious correlation causes machine learning models to fail on certain groups of samples, even when achieving high accuracy on average group (Ribeiro et al., 2016; Beery et al., 2018; Hashimoto et al., 2018; Duchi et al., 2019). Deep networks trained via standard empirical risk minimization (ERM) may be biased by the spurious attributes/features such as the image backgrounds (Beery et al., 2018; Geirhos et al., 2020), which admit different correlations with the true labels for the different groups of data. As a result, the model typically performs well on the majority group of data that aligns with the spurious attributes but performs poorly on certain groups of data that have no or even opposite correlations with spurious attributes. This critical issue is widespread in many fields, such as medical AI, facial recognition and sentiment analysis (Blodgett et al., 2016; Tatman, 2017; Gururangan et al., 2018; Badgeley et al., 2019; Sagawa et al., 2019).
@@
There have emerged a series of previous works that aim to develop more robust training methods to improve the worst-group performance when group information is available. For instance, one can focus on particularly training over the worst-group risk (Sagawa et al., 2019) via group distributional robust optimization (GroupDRO). Additionally, the spurious correlation can also be eliminated by balancing the risk in different groups, which can be achieved via feature reweighting (Kirichenko et al., 2022) or selective Mixup (Yao et al., 2022). These methods have been demonstrated to be effective to mitigate the spurious correlations, leading to substantial improvements over ERM.
@@
However, group labeling is expensive, labor-intensive, and human-biased. It is more important and practical to develop robust algorithms for improving worst-group accuracy when the group information is unavailable (Liu et al., 2021; Zhang et al., 2022). In such a scenario, a line of recent works (Sohoni et al., 2020; Nam et al., 2020; Ahmed et al., 2020; Liu et al., 2021; Zhang et al., 2022; Chen et al., 2023) propose to first infer the data groups, and then seek to train a robust model according to the estimated group information. For instance, JTT (Liu et al., 2021) identifies the data being misclassified by the ERM as the minority group, and trains a more robust model via upweighting the examples in minority groups; CnC (Zhang et al., 2022) infers pseudo group labels via ERM and develops a contrastive learning method across groups to train the final model.
@@
The primary focus of the aforementioned research is on invariant learning, i.e., leveraging the group information obtained from ERM to learn the invariant attributes. However, the accuracy of group label inference has been overlooked, making the existing group inference-based methods perform significantly worse than those using oracle group labels. Very recently, some group inference methods have been developed as alternatives to ERM, such as EIIL (Creager et al., 2021), SSA (Nam et al., 2022), ZIN (Lin et al., 2022) and DISC (Wu et al., 2023). However, they may either struggle with complex spurious correlations (e.g., EIIL) (Lin et al., 2022), or require additional information (e.g., human inspection, few spurious annotations, etc) that is related to the group labels (e.g., SSA, ZIN, and DISC). Therefore, noticing the performance shortcomings and applicability limitations in existing group inference methods, we raise an important yet challenging question:
Can we develop a more accurate group inference method to mitigate spurious correlations without relying on any additional information?
@@
In this work, we propose GIC (Group Inference via data Comparison), a novel, practical and accurate method to infer group labels, which can be then seamlessly incorporated into a variety of oracle group label-based robust learning methods to improve the worst-group performance. In particular, GIC adopts a (unlabeled) comparison dataset, which has (slightly) different group distribution compared to the training dataset, and infers the group information in a contrastive manner. More specifically, GIC trains an alternative classifier to assign spurious attribute labels to data points solely on their spurious attributes, where the training objective is designed to encourage discovering the distribution discrepancy between the comparison and training dataset while maintaining the high correlation between the spurious attribute label and true label in the training dataset. Notably, unlike previous methods that may rely on (partial) group labels or human inspections, obtaining the comparison data does not necessitate any assumptions. It does not need to be labeled and can be obtained from various sources, such as a validation set with true validation labels, a completely unlabeled subset sampled from the test data, or even directly resampled from the training dataset in a non-uniform manner. We highlight the main contributions as follows:
@@
    We show that the proposed GIC can seamlessly integrate with multiple invariant learning algorithms to improve the worst-group accuracy. In Section 4.3 and 4.4, GIC is successfully combined with Mixup (Yao et al., 2022), GroupDRO (Creager et al., 2021), Upsample (Liu et al., 2021), and Subsample (Kirichenko et al., 2022). It can be seen that GIC consistently outperform baselines in terms of the worst-group accuracy, for all candidate invariant learning algorithms. More importantly, when integrated with Mixup, our model improves over the state-of-the-art for nearly all tasks when the group label is unavailable. Additionally, the average and worst-group accuracy of our model can almost match that of using oracle group labels directly. This further justifies the effectiveness of our group inference method.
@@ 2 Related Work
Recent research shows that the traditional ERM can learn both spurious and invariant features (Kirichenko et al., 2022; Izmailov et al., 2022; Rosenfeld et al., 2022; Chen et al., 2023). However, with strong spurious correlations, ERM tends to prioritize learning the spurious features (Chen et al., 2023), which hinders its ability to generalize on data where these spurious correlations are absent. To tackle this issue, invariant feature learning, referred to as invariant learning for simplicity, has been introduced to learn invariant representations. A notable approach to invariant learning is the use of group robustness methods.
When group labels are available, some classical group robustness methods, such as GroupDRO (Sagawa et al., 2019), attempt to minimize the worst-group loss instead of the average loss using oracle group labels. Other methods aim to achieve invariant learning by balancing the majority and minority groups, such as reweighting (Sagawa et al., 2020), regularization (Cao et al., 2019), and downsampling (Kirichenko et al., 2022). Additionally, approaches like semi-supervised learning (Nam et al., 2022; Sohoni et al., 2021) or Mixup (Yao et al., 2022), which selectively combine samples with matching labels but differing spurious attributes or matching spurious attributes but differing labels, are also be considered to improve the the worst-group accuracy. Some methods attempt to infer group labels by training a simple ERM (Blodgett et al., 2016; Tatman, 2017; Gururangan et al., 2018; Badgeley et al., 2019; Sagawa et al., 2019), maximizing the GroupDRO loss (Creager et al., 2021), or even introducing human knowledge (Lin et al., 2022; Wu et al., 2023) when group information is unknown. However, these group inferred methods have performance gaps compared to group annotation utilized methods and may not be applicable when prior information is unavailable. In this work, we focus on group robustness without relying on any group labels or human-provided information.
We point out that, although we aims to learn invariant features to train more robust and generalizable models, similar to classical domain adaptation (Blanchard et al., 2011; Muandet et al., 2013) methods, such as UDA (Ganin & Lempitsky, 2015) and DANN (Ganin et al., 2016), our setting differs from them which divide the training data into source and target domains and use the source and target features, along with the source labels, to transfer knowledge to the specific target domain (Zhang et al., 2022). In contrast, we do not have a natural source and target domains and strive to enhance the accuracy of specific groups affected by spurious correlations, without prior knowledge of training sample domains or spurious attributes.
@@ 3.1 Problem Setup
Consider the training dataset $\mathcal{D}$, which comprises $n$ data point-label pairs, where each data point-label pair belongs to some group $g\in\mathcal{G}$, denoted as $\mathcal{D}=\{(\mathbf{x}_{i},y_{i},g_{i})\}_{i=1}^{n}$. Following previous work (Liu et al., 2021; Zhang et al., 2022; Yao et al., 2022; Yang et al., 2023), we consider each group $g=(y,a)$ to be jointly defined by the label $y\in\mathcal{Y}$ and unobserved spurious attributes $a\in\mathcal{A}$ (i.e., $\mathcal{G}=\mathcal{Y}\times\mathcal{A}$), where the spurious attribute $a$ is spuriously correlated with the label $y$ (e.g., the image background and the image label). We denote $y_{s}$ as the spurious attribute label of $a$ (e.g., the label of the image background, which can be “land” versus “water” or “red” versus “green”). The groups $g$ in dataset $\mathcal{D}$ often encounter the imbalance issue, where ERM training tends to mostly depend on majority groups, thus may memorize the spurious correlation contributed by these groups. However, such spurious correlations are often absent or even oppositely appearing in minority groups. Therefore, our ultimate objective is to train a robust model $f_{\mathrm{robust}}(\theta)$, parameterized by $\theta\in{\Theta}$, capable of classifying an input $\mathbf{x}_{i}$ to a label $y_{i}$ regardless of whether it belongs to the majority or minority group. This is referred to as the worst-group accuracy defined as follows:
${\max_{g\in\mathcal{G}}\mathbb{E}[\mathds{1}[f_{\theta}(\mathbf{x})\neq y]|g]},$ (1)
@@ 3.2 GIC: Group Inference via data Comparison
$\hat{y}_{s,\mathbf{w}}=f_{\mathrm{GIC}}(\mathbf{z};\mathbf{w}),$ (3)
where $\mathbf{w}$ is weights of GIC model $f_{\mathrm{GIC}}$. Then using the predicted $\hat{y}_{s,\mathbf{w}}$, GIC partitions the data into different groups, i.e., $\hat{g}=(y,\hat{y}_{s,\mathbf{w}})$. The inferred group $\hat{g}$ can be integrated into downstream invariant learning methods, aiding in training the robust model $f_{\mathrm{robust}}$.
Comparison data. Before formally introducing details of GIC, we first introduce the concept of comparison data. We assume access to a dataset $\mathcal{C}$, where its group distribution (slightly) differs from the training data $\mathcal{D}$, e.g., the group distribution of $\mathcal{D}$ is $(g_{1},g_{2})=(0.1,0.9)$, while the group distribution of $\mathcal{C}$ is $(g_{1},g_{2})=(0.2,0.8)$. We refer to $\mathcal{C}$ as comparison data, which can have true labels or be unlabeled. We remark that the comparison data is easy to obtain, which will be thoroughly discussed in Section 3.4.
Based on the definition of spurious attribute labels $y_{s}$, an ideal spurious attribute prediction $\hat{y}_{s,\mathbf{w}}$ should exhibit the following two properties: (1) It should be highly correlated with the true label $y$. It is this high correlation between the spurious attribute label and the true label that biases the ERM-based model training. (2) The correlation between $\hat{y}_{s,\mathbf{w}}$ and $y$ varies across different datasets. Unlike the invariant attribute label, which is always equivalent to the true label, the correlation between the spurious attribute label and the true label is spurious, changing with the spurious attribute distribution across different datasets. Consider an extreme case, the spurious background attribute for both waterbirds and landbirds in the comparison data accounts for 50%, while in the training set, the spurious attribute label is completely equivalent to the true label (e.g., $y_{s}=y$). The equivalence correlation that exists in the training data does not hold in the comparison data.
Based on the mentioned two properties, we design the optimization objective of GIC, which consists of two terms: (1) Correlation Term, used to describe the high correlation between $\hat{y}_{s,\mathbf{w}}$ and $y$ in the training data; (2) Spurious Term, used to describe the discrepancy in the correlation between the training and comparison data, aiming to emphasize the correlation is spurious rather than invariant. By jointly optimizing the above objectives, GIC is encouraged to train a spurious attribute classifier that yields the spurious attribute prediction $\hat{y}_{s,\mathbf{w}}$, as shown in Equation (3).
Correlation Term. To encourage the high correlation between $y$ and $\hat{y}_{s,\mathbf{w}}$ in the training set, we consider the following optimization objective:
@@
where $\hat{y}^{tr}_{s,\mathbf{w}}=f_{\mathrm{GIC}}(\mathbf{z}^{tr};\mathbf{w})$ is the predicted spurious attribute label on training data. Here, mutual information is considered due to its widespread usage to measure the correlation or dependency between random variables (Li & Jurafsky, 2016; Kong et al., 2019; Pan et al., 2020; Su et al., 2023). We aim to maximize the mutual information between $y_{s}$ and $y$, encouraging the predicted spurious attribute label $y_{s}$ from GIC to be highly correlated with the true label $y$.
Solely satisfying Equation (4) is not enough as invariant attribute information contained in feature representations (Kirichenko et al., 2022; Chen et al., 2023) can also exhibit high correlation with true labels. To aid GIC in learning spurious attributes instead of invariant ones, we introduce an additional spurious term.
Spurious Term. We use conditional probability to describe the correlation between $\hat{y}_{s,\mathbf{w}}$ and $y$, i.e., $\mathbb{P}(y|\hat{y}_{s,\mathbf{w}})$. Then let $\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s})$ and $\mathbb{P}(y^{c}|\hat{y}^{c}_{s})$ be the conditional distributions in the training and comparison dataset respectively, we will characterize their discrepancy. Since we are comparing distributions with the same support, we opt for the simplest and widely used choice (Ahmed et al., 2020; Lee et al., 2022), the KL-divergence. Then, we maximize the following objective to encourage GIC to learn spurious attributes:
@@
Note that a more formal expression for Equation (5) should use different distribution notations and the same variable names, i.e.,
$\max_{\mathbf{w}}\mathrm{KL}(\mathbb{P}_{\mathrm{tr}}(y|\hat{y}_{s,\mathbf{w}})||\mathbb{P}_{\mathrm{c}}(y|\hat{y}_{s,\mathbf{w}}))$ (6)
where $\mathbb{P}_{tr}(y|\hat{y}_{s,\mathbf{w}}):=\mathbb{P}(y^{\mathrm{tr}}|\hat{y}^{tr}_{s,\mathbf{w}})$ and $\mathbb{P}_{\mathrm{c}}(y|\hat{y}_{s,\mathbf{w}}):=\mathbb{P}(y^{c}|\hat{y}^{c}_{s,\mathbf{w}})$. Here, distinct variable names, such as $\hat{y}^{tr}_{s,\mathbf{w}}$ and $\hat{y}^{c}_{s,\mathbf{w}}$, are utilized to indicate their origins from different (training or comparison) datasets, thereby emphasizing GIC can achieve improved group inference via data comparison.
The essence of maximizing Equation (5) is to encourage GIC to learn spurious attributes by violating the invariant learning principle (Creager et al., 2021). If $\hat{y}_{s,\mathbf{w}}$ is inferred based on invariant attributes rather than spurious ones, then $\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})$ and $\mathbb{P}(y^{c}|\hat{y}^{c}_{s,\mathbf{w}})$ will be identical and then
@@
Conversely, if $\hat{y}_{s,\mathbf{w}}$ is based on spurious attributes rather than invariant attributes, the inequality
@@
holds, where the equality holds only when the training and comparison data share the same group distribution. Therefore, maximizing Equation (5) entails encouraging GIC to prioritize learning spurious attributes over invariant ones.
@@
  Obtain $\mathbf{z}^{tr}=\Phi(\mathbf{x}^{tr}),\mathbf{z}^{c}=\Phi(\mathbf{x}^{c})$ where $\mathbf{x}^{tr}\in\mathcal{D},\mathbf{x}^{c}\in\mathcal{C}$.
@@
  Initialize the parameters $\mathbf{w}$ for spurious attribute classifier $f_{\mathrm{GIC}}$.
  for epoch $1$ to $K$ do
     if the true label of $\mathcal{C}$ is available then
        Optimizing Equation (11) to update $\mathbf{w}$.
     else
        Optimizing Equation (12) to update $\mathbf{w}$.
     end if
  end for
  Infer spurious attribute labels $\hat{y}^{tr}_{s,\mathbf{w}}=f_{\mathrm{GIC}}(\mathbf{z}^{tr};\mathbf{w})$.
  Return: Pseudo group labels $\hat{g}=(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$.
Overall Objective. By combining Equation (4) and (5), we derive the overall objective of GIC as follows:
@@
where $\gamma\geq 0$ is a weighting parameter used to balance Correlation Term and Spurious Term.
Figure 1: Decision boundary visualization. $f_{\mathrm{ERM}}$ underperforms $f_{\mathrm{GIC}}$ in recognizing spurious attributes and $f_{\mathrm{robust}}$ in identifying invariant attributes. Classes 00 and $1$ are represented by colors (red and blue), respectively, with shapes marking spurious attributes.
Figure 2: Evaluation of group label inference. Compared to baseline methods such as ERM and EIIL, GIC significantly improves the recall for minority group label inference while maintaining a relatively high precision.
However, the overall objective of GIC faces certain practical issues. Firstly, the mutual information term is difficult to accurately estimate (Paninski, 2003; Belghazi et al., 2018). Secondly, Equation (9) cannot handle the situation where the comparison data is unlabeled, which limits the applicability of GIC in various scenarios.
We first replace the mutual information $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$ with the cross-entropy $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$ to achieve accurate estimation. A detailed proof in Appendix A.1 demonstrates that $-H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$ is, in fact, a lower bound of $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$. Therefore, maximizing $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$ can be achieved by minimizing $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$. This replacement aligns with intuition because maximizing mutual information between $y^{tr}$ and $\hat{y}^{tr}_{s,\mathbf{w}}$ essentially encourages a closer alignment of their distributions, which is consistent with the objective of minimizing cross-entropy $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$.
We then extend GIC to the case where the true label $y^{c}$ of the comparison data are not available. We substitute the true labels $y^{c}$ with the comparison data’s feature representation $\mathbf{z}^{c}$, which strongly associates with $y^{c}$ and is always accessible. We present the following theorem: